# RPCBF: Constructing Safety Filters Robust to Model Error and Disturbances via Policy Control Barrier Functions

### v2

@@ III-C Time Discretization of Policy Control Barrier Functions
Another challenge lies in how to compute the maximum in (10). The states $\mathbf{x}_{t}^{\pi}$ can be solved numerically using an ODE solver, resulting in a time-discretized state trajectory. It is tempting to then consider taking the maximum $h$ over this trajectory, i.e., for time discretization $\Delta t$,
@@
Figure 2: Value Function Gradient Error for Discrete-Time Double Integrator. We highlight the discretized trajectory (top) and corresponding gradient error (bottom) for three different choices of $\Delta t$ (yellow, purple, grey). The gradient of the naive discrete-time approximation has large errors and varies with the choice of $\Delta t$. Taking the maximum of the cubic spline leads to much smaller errors.
Figure 3: Summary of RPCBF Algorithm. Given a policy $\pi$, we sample disturbance trajectories, then compute the maximum $h$ with cubic splines to obtain $V^{h,\pi}$ and $\nabla V^{h,\pi}$ (using autodiff). This is used in a CBF-QP to obtain a robust safety filter.
@@
Improved Time-Discretization using Cubic Splines. To reduce the error in the time-discretized value function approximation (13), we propose to approximate $h(\mathbf{x}_{t}^{\pi})$ by fitting a cubic spline to the points $\{h(\mathbf{x}_{k\Delta t}^{\pi})\}_{k=0}^{H-1}$. The $\max$ over the cubic spline can then be computed in closed-form by solving the roots of a quadratic to yield a better approximation of $\sup_{0\leq t<T}h(\mathbf{x}_{t}^{\pi})$ than the naive maximization (13). Intuitively, this resolves the gradient error due to integer-valued $\tau$ from the previous example because the maximum of the cubic spline can now happen between timesteps. Applying this method to the previous example of the DI results in a significantly smaller gradient error (fig. 2).
@@ III-D Robust Extension of PCBFs
In this subsection, we now consider a robust extension of PCBF to handle disturbances. Defining the robust value function equivalent of (6) as
@@
Solving for robust controls that satisfy (18) renders the zero sublevel set robust forward-invariant [17].
However, deriving the worst-case disturbance is generally intractable because it requires evaluating all possible disturbance trajectories. Instead, we propose to only consider $N$ disturbance trajectories and take the worst-case out of the $N$ samples, resulting in the following practical Robust Policy Control Barrier Function (RPCBF) approximation.
$\dot{\mathbf{x}}^{i}$ $=f(\mathbf{x}^{i},\mathbf{d}^{i})+g(\mathbf{x}^{i},\mathbf{d}^{i})\mathbf{u}^{i}.$ (20)
We summarize our approach in algorithm 1 and fig. 3. Note that algorithm 1 must be executed once per control loop to obtain the value $V^{h,\pi}_{T,N}$ and gradient $\nabla V^{h,\pi}_{T,N}$ for the current state for use in the CBF-QP (CBF-QP). Different approaches can be implemented to perform informed sampling of disturbances. For bounded disturbances, the worst-case scenario often occurs at the vertices of the disturbance set (e.g., for disturbance-affine dynamics). Consequently, we choose to sample from a mixture of the uniform distribution $\mathcal{U}(\mathbf{d}_{\mathrm{min}},\mathbf{d}_{\mathrm{max}})$ and the uniform distribution over the vertices. This can be extended to better optimizers to approximate the worst-case samples which we leave as future work.
While the finite-sample approximation does not guarantee robustness to any disturbance, it does ensure robustness to the specific sampled disturbances within the finite horizon. As the number of informed samples approaches infinity, the approximation increasingly captures the true worst-case scenarios. For the simulation and hardware experiments below, we use PCBF and RPCBF to refer to their time-discretized finite-horizon and finite-sample approximations as described in this section.
@@
1:Input: Initial State $\mathbf{x}_{0}$, Policy $\pi$, Constraint function $h$, Horizon $T=H\Delta t$ , Number of disturbance samples $N$
@@
8:Compute the gradient $\nabla V^{h,\pi}_{T,N}(\mathbf{x}_{0})$ using autodiff
@@ IV Simulation Experiments
To study the performance of PCBF and RPCBF, we perform a series of simulation experiments on high relative degree systems under box control constraints.
Baselines. We compare against the following SFs that also do not use NNs in their approach.
@@
  • Approximate Nominal MPC-based Predictive Safety Filter (MPC) [44]: A trajectory optimization problem is solved, imposing the safety constraints while penalizing deviations from the nominal policy. We consider the undisturbed system and do not assume access to a known robust forward-invariant set and hence do not impose this terminal constraint.
Systems. We consider three systems: a DI, a Segway, and AutoRally [45], an 1/5 autonomous vehicle. On the DI, safety is defined via position bounds ($|p|\leq 1$), while the segway asks for the handlebar to stay upright and for position bounds to be maintained ($|\theta|\leq 0.3\pi$, $|p|\leq 2)$, $\Delta t=0.1$. For AutoRally, a crash occurs when the car hits the track boundary and comes to a complete stop, while a collision is defined as the car making contact with the track boundary without coming to a halt. For the DI and the Segway, we assume an unknown but bounded mass and treat the unknown mass as a disturbance to the system. For the AutoRally, we assume additive truncated Gaussian noise. During testing, we consider a constant zero-control nominal policy for the DI, a maximum-acceleration control nominal policy for the segway, and use model predictive path integral (MPPI) [46] as the nominal policy for AutoRally. Note that the nominal policy $\pi_{\mathrm{nom}}$ differs from the policy $\pi$ used to derive the PCBF.
Figure 4: Filter Boundary and Safe Region for PCBFs with varying Horizon $\boldsymbol{T}$ on the Segway. We plot the states where the nominal policy can influence the output of the safety filter (Filter Boundary) and the initial states from which the safety filter can preserve safety over a horizon $\bar{T}=30s$ (Safe Region). A trajectory from an initial state inside the filter boundary marked by a black dot ($\bullet$) is displayed. If the horizon is too short, the filter boundary is overapproximated, resulting in an unsafe trajectory.
IV-A Influence of Horizon Length on Segway
While the infinite-horizon policy value function is a CBF, we consider a finite-horizon approximation. The performance of the SF is thus dependent on the chosen horizon length. To gain a better intuition, we evaluate how different horizon lengths affect the performance of the PCBF-SF. Here, we consider the undisturbed Segway.
We visualize the results in fig. 4, plotting the state space from where $\pi_{\mathrm{nom}}$ can influence the output of the SF (Filter Boundary) and from where the SF preserves safety (Safe Region). For CBF-based filters, the Filter Boundary is defined by the zero level set of the CBF. The Safe Region is determined for a $\pi_{\mathrm{nom}}$ by solving (CBF-QP) and rolling out the system over a horizon $\bar{T}=30s$. For a short PCBF horizon, i.e., $T=5s$, the true Control Invariant (CI) set is overapproximated. Consequently, the SF fails to preserve safety, as illustrated by the resulting unsafe example trajectory. In contrast, a longer horizon of $T=20s$ provides a much closer approximation of the true CI set. This is evident when comparing it to an even longer horizon, such as $T=30s$, which does not result in a visibly larger safe region, indicating that $T=20s$ is already sufficient to capture the safety requirements accurately.
@@ IV-A Influence of Horizon Length on Segway
Figure 5: Comparison of Filter Boundary and Safe Region on the Double Integrator (a) and Segway (b). The true unsafe region for the undisturbed system is shaded in gray for the double integrator. PCBF-based methods use a horizon $T$ and $N$ samples to derive the policy value function. The Safe Region is determined for a horizon $\bar{T}$. Trajectories from selected initial states for $\bar{N}=25$ sampled disturbance trajectories are displayed. Red dotted lines and green solid lines indicate unsafe and safe trajectories, respectively. The nominal trajectory is shown in black.
Figure 6: Robust Safety Evaluation. We plot the zero level set of the CBFs. Green dots indicate safe rollouts for all $\bar{N}=25$ sampled disturbance trajectories, while red crosses indicate a failure in at least one trajectory. RPCBF achieved safety for all states within its zero level set for the sampled disturbances.
@@ IV-B Behavior on Disturbed Double Integrator and Segway
Next, we explore the robustness of the different SFs. We again examine the filter boundary and safe region, this time they are derived for one sampled disturbance/ disturbance trajectory per state and are illustrated in fig. 5. We visualize rolled-out trajectories for $\bar{N}$ sampled disturbances (uniformly sampled and on the vertices) from initial states within the filter boundary of the considered method. On the DI, only RPCBF achieved safety for all $\bar{N}$ sampled disturbances. Since RPCBF accounts for the worst-case scenario among the sampled disturbance trajectories, the filter boundary is more conservative. On the segway, MPC violated safety constraints in all cases and hence has an empty filter boundary and safe region. Only RPCBF achieves safe trajectories for all considered samples. Furthermore, we evaluate the PCBF-based SFs on states uniformly distributed across the state space for $\bar{N}$ sampled disturbances. The results are displayed in fig. 6. The RPCBF-based SF achieves safety for all evaluated states within its zero level set.
@@ IV-C Simulations on AutoRally
Finally, to assess the safety improvements brought about by the proposed PCBF and RPCBF, we integrate the HOCBF and the proposed methods with Shield-MPPI [45, 47], and evaluate their performance on AutoRally. The fig. 7 shows that Shield-MPPI using RPCBF generates the safest trajectories, while other controllers generate trajectories that collide and crash more often. The statistics of the safety performance of the controllers are shown in table I.
Figure 7: Trajectory Comparisons on AutoRally. We visualize the trajectories for each controller, where SMPPI-RPCBF (in blue) leads to the tightest spread of states inside the track.
TABLE I: Collision and Crash Rate on AutoRally. The standard MPPI leads to most collisions and crashes. While Shield-MPPI with HOCBF and PCBF significantly improves safety, Shield-MPPI using RPCBF achieves the lowest rate of collision and crashes.
@@ V Hardware Experiments
Figure 8: Hardware Traces. RPCBF maintains safety while following the unsafe nominal trajectory despite error between the modeled and true Crazyflie dynamics. On the other hand, PCBF assumes the modeled dynamics are perfect and collides with the obstacle.
Figure 9: Safety with different values of $\alpha$. We plot the minimum distance to the obstacle, over $6$ random nominal trajectories. Although PCBF is safe when $\alpha$ is sufficiently small, this requires fine-tuning and is difficult to know beforehand. RPCBF is safe for all values of $\alpha$ we tested.
@@
We first run the PCBF and RPCBF controllers with $\alpha=5$ on $6$ different nominal trajectories and plot the results from $3$ of the random trajectories in fig. 8. The RPCBF maintains safety in all cases, while the PCBF collides in all cases.
We next vary the choice of the class-$\kappa$ function $\alpha$ and plot the results in fig. 9. While the non-robust PCBF does not collide with the obstacle when $\alpha$ is sufficiently small, this requires fine-tuning and is difficult to know beforehand. On the other hand, RPCBF is safe for all values of $\alpha$ we tested, allowing $\alpha$ to be used as a parameter that controls the behavior without also simultaneously affecting safety.
@@ VI Discussion and Conclusion
In this work, we have proposed the Robust Policy Control Barrier Function (RPCBF), a method of constructing SFs robust to disturbances via robust CBFs. Simulation experiments show that RPCBF yields improved safety and more accurate estimation of the robust control-invariant set compared to existing methods. Hardware experiments on the quadcopter show that considering model errors is crucial to maintaining safety.
Future directions include performing a more thorough analysis of the safety guarantees of RPCBFs. In particular, although we have provided a simple theorem proving that a long enough horizon is sufficient, this does not account for time-discretization of the rollouts. Also, as noted in section IV-A, the forward-invariant set of the RPCBF is sometimes larger than the zero sublevel set. Whether rigorous statements can be made about this case is another interesting future direction to explore.

### v3

@@ III-C Time Discretization of Policy Control Barrier Functions
Another challenge lies in how to compute the maximum in (10). The states $\mathbf{x}_{t}^{\pi}$ can be solved numerically using an ordinary DE solver, resulting in a time-discretized state trajectory. It is tempting to then consider taking the maximum $h$ over this trajectory, i.e., for time discretization $\Delta t$,
@@
Improved Time-Discretization using Cubic Splines. To reduce the error in the time-discretized value function approximation (13), we propose to approximate $h(\mathbf{x}_{t}^{\pi})$ by fitting a cubic spline to the points $\{h(\mathbf{x}_{k\Delta t}^{\pi})\}_{k=0}^{H-1}$. The $\max$ over the cubic spline can then be computed in closed-form by solving the roots of a quadratic to yield a better approximation of $\sup_{0\leq t<T}h(\mathbf{x}_{t}^{\pi})$ than the naive maximization (13). Intuitively, this resolves the gradient error due to integer-valued $\tau$ from the previous example because the maximum of the cubic spline can now happen between timesteps. We can also formally quantify the error in both the cubic spline value and its gradient. Let $\tilde{h}:[0,\infty)\to\mathbb{R}$ denote the cubic spline approximation of $h$ as a function of time, and let $\tilde{V}^{h,\pi}_{\infty,\Delta t}(\mathbf{x}_{0})\coloneqq\sup_{t\geq 0}\tilde{h}(t)$. Using [15, Chapter 5], we obtain the error bounds
$\left\lVert V^{h,\pi}_{\infty}-\tilde{V}^{h,\pi}_{\infty,\Delta t}\right\rVert$ $\leq\frac{1}{16}\Delta t^{4}\max_{t\geq 0}\left\lVert\frac{\mathop{}\!\mathrm{d}^{4}}{\mathop{}\!\mathrm{d}t}h(\mathbf{x}_{t})\right\rVert,$ (17a)
$\left\lVert\nabla V^{h,\pi}_{\infty}-\nabla\tilde{V}^{h,\pi}_{\infty,\Delta t}\right\rVert$ $\leq\frac{1}{24}\Delta t^{3}\max_{t\geq 0}\left\lVert\frac{\mathop{}\!\mathrm{d}^{4}}{\mathop{}\!\mathrm{d}t}h(\mathbf{x}_{t})\right\rVert.$ (17b)
In the previous example of the DI, since $h$ is exactly quadratic, applying cubic splines results in zero gradient error (fig. 2). If $\mathop{}\!\mathrm{d}^{4}/\mathop{}\!\mathrm{d}t^{4}\;h(\mathbf{x}_{t})$ can be bounded, the bounds (17) can then be used to suitably modify (CBF-QP) to guarantee safety. A larger $\Delta t$ will therefore result in a more conservative SF.
Figure 3: Summary of RPCBF Algorithm. Given a policy $\pi$, we sample disturbance trajectories, then compute the maximum $h$ with cubic splines to obtain $V^{h,\pi}$ and $\nabla V^{h,\pi}$ (using automatic differentiation). This is used in (CBF-QP) to obtain a robust SF.
@@ III-D Robust Extension of PCBFs
We extend PCBF to handle disturbances by defining the robust value function equivalent of (6) as
@@
1:Input: Initial State $\mathbf{x}_{0}$, Policy $\pi$, Constraint function $h$, Horizon $T=H\Delta t$ , Number of disturbance samples NN
@@
8:Compute the gradient $\nabla V^{h,\pi}_{T,N}(\mathbf{x}_{0})$ using automatic differentiation
Solving for robust controls that satisfy (19) renders the zero sublevel set robust forward-invariant [18]. However, deriving the worst-case disturbance is generally intractable because it requires evaluating all possible disturbance trajectories. Instead, we propose to only consider $N$ disturbance trajectories and take the worst-case out of the $N$ samples, resulting in the following RPCBF approximation:
$\dot{\mathbf{x}}^{i}_{t}$ $=f(\mathbf{x}^{i}_{t},\mathbf{d}^{i}_{t})+g(\mathbf{x}^{i}_{t},\mathbf{d}^{i}_{{\color[rgb]{0,0,0}t}})\mathbf{u}^{i}_{t}.$ (21)
We summarize our approach in algorithm 1 and fig. 3. Note that algorithm 1 must be executed once per control loop to obtain the value $V^{h,\pi}_{T,N}$ and gradient $\nabla V^{h,\pi}_{T,N}$ for the current state for use in (CBF-QP). Different approaches can be implemented to perform informed sampling of disturbances. For bounded disturbances, the worst-case scenario often occurs at the vertices of the disturbance set (e.g., for disturbance-affine dynamics). Consequently, we choose to sample from a mixture of the uniform distribution $\mathcal{U}(\mathbf{d}_{\mathrm{min}},\mathbf{d}_{\mathrm{max}})$ and the uniform distribution over the vertices. Using better optimizers to approximate the worst-case samples will be left to future work.
While the finite-sample approximation does not guarantee robustness to any disturbance, it does ensure robustness to the specific sampled disturbances within the finite horizon. As the number of informed samples approaches infinity, the approximation increasingly captures the true worst-case scenarios. Theoretical guarantees for this sampling-based approach could be established using random set theory [28], as demonstrated in [21]. Furthermore, statistical risk measures could be used instead of the worst-case out of $N$ samples. For instance, using [3] who bound the risk measure evaluation of a random variable whose distribution is unknown. However, we leave this for future work. For the simulation and hardware experiments below, we use PCBF and RPCBF to refer to their time-discretized finite-horizon and finite-sample approximations as described in this section.
@@ IV Simulation Experiments
We evaluate the performance of (R)PCBF in simulation for high relative degree systems with box control constraints.
Baselines. We compare against the following SFs, which similarly do not incorporate NNs in their approach.
@@
  • Approximate Nominal MPC-based PSF (MPC) [42]: A trajectory optimization problem is solved, imposing the safety constraints while penalizing deviations from the nominal policy. We consider the undisturbed system without assuming access to a known robust forward-invariant set, thus not imposing a terminal constraint.
Systems. We consider four systems: a DI, a Segway, an F-16 fighter jet (ground collision avoidance problem) [16, 38], and AutoRally [48], a 1/5 autonomous vehicle. On the DI, we consider position bounds ($|p|\leq 1$), while the Segway asks for an upright handlebar and considers position bounds ($|\theta|\leq 0.3\pi$, $|p|\leq 2)$, $\Delta t=0.1$. For the F-16, safety is defined as box constraints on states like altitude. Since this system is not control-affine in the throttle, we leave the throttle as the output of a P controller, resulting in a $16$-dimensional state space and a $3$-dimensional control space. In AutoRally, a crash occurs when the car stops after hitting the track boundary, while a collision involves contact without stopping. For each system, we define $J$ continuously differentiable constraint functions $h_{j}$ tailored to the problem at hand. For example, for the DI system we set $h_{0}=p-1$ and $h_{1}=-(p+1)$. From these we derive $J$ corresponding CBFs, which yield to $J$ constraints in (CBF-QP). For the DI and the Segway, we assume unknown but bounded time-varying disturbances on the mass, for the F-16 unknown but bounded matched disturbances ($d=1$), and for the AutoRally additive truncated Gaussian noise. Keep in mind that our approach does not require designing/learning a new CBF for different systems, disturbance assumptions, or input constraints, but simply requires swapping the dynamics, specifying the disturbance and constraints. During testing, we consider a constant zero-control nominal policy for the DI, maximum acceleration for the Segway, a PID controller for the F-16, and Model Predictive Path Integral (MPPI) control [50] for AutoRally.
Figure 4: Filter Boundary and Safe Region for PCBF-SFs with varying Horizon $\boldsymbol{T}$ on the Undisturbed Segway. We plot where the nominal policy affects the SF’s output (Filter Boundary) and the states where the SF ensures safety over $\bar{T}=30s$ (Safe Region). Trajectories from an initial state within the filter boundary (black dot, $\bullet$) are color-coded by horizon length of the PCBF-SF. A too-short horizon overapproximates the filter boundary, causing unsafe trajectories.
@@
Figure 5: Comparison of Filter Boundary and Safe Region on DI (a) and Segway (b). The true unsafe region for the undisturbed DI is shaded in gray. (R)PCBFs use horizon $T$ and NN samples to derive the value function. The safe region is determined for $\bar{T}$. Trajectories from selected initial states are shown for $\bar{N}=25$ sampled $\mathbf{d}$ trajectories. Red dotted and green solid lines indicate unsafe and safe trajectories, respectively, with the nominal trajectory in black.
IV-A Influence of Horizon Length on Undisturbed Segway
While the infinite-horizon policy value function is a CBF, we use a finite-horizon approximation, making the SF performance horizon-dependent. To illustrate this, we assess the impact of different horizon lengths on the PCBF-SF, see fig. 4. We plot the state space from where $\pi_{\mathrm{nom}}$ can influence the output of the SF (Filter Boundary) and from where the SF preserves safety (Safe Region). For CBF-based filters, the filter boundary is defined by the CBF’s zero level set. The safe region is determined for a $\pi_{\mathrm{nom}}$ by solving (CBF-QP) and rolling out the system over a horizon $\bar{T}=30s$. For a short PCBF horizon, i.e., $T=5s$, the true CI set is overapproximated. Consequently, the SF fails to preserve safety, as illustrated by the resulting unsafe example trajectory. In contrast, a longer horizon of $T=10s$ provides a much closer approximation of the true CI set. This is evident when comparing it to an even longer horizon, such as $T=30s$, which does not result in a visibly smaller safe region, indicating that $T=10s$ is already sufficient.
@@ IV-B Behavior on Disturbed Double Integrator and Segway
We explore the robustness of different SFs, examining the filter boundary and safe region, derived for one sampled disturbance trajectory per state, as shown in fig. 5. We visualize rolled-out trajectories for $\bar{N}$ sampled disturbance trajectories (uniformly sampled and on the vertices) from selected initial states within the filter boundary. On the DI, only the RPCBF-SF achieves safety for all $\bar{N}$ sampled disturbance trajectories. Since the RPCBF accounts for the worst-case among the $N$ sampled disturbances, the filter boundary is more conservative. On the Segway, MPC violates the safety constraints in all cases and hence has an empty filter boundary and safe region. Only the RPCBF-SF achieves safe trajectories for all considered samples. Next, we evaluate the (R)PCBF-SFs at uniformly distributed initial states, see fig. 6. The RPCBF-SF achieves safety for all evaluated states within its zero-level set and the sampled disturbances, while the PCBF overapproximates the safe set.
Figure 6: Robust Safety Evaluation. We plot the zero level set of the (R)PCBFs. Green dots indicate safe rollouts for all $\bar{N}=25$ sampled disturbance trajectories, while red crosses indicate a failure in at least one trajectory. The RPCBF-SF achieved safety for all states within its zero level set for the sampled disturbance trajectories.
@@ IV-C Simulations on AutoRally
Finally, to assess the safety improvements brought about by the proposed (R)PCBF, we integrate the HOCBF and the proposed methods with Shield-MPPI (SMPPI) [48, 49], and test them on AutoRally. Figure 7 shows that SMPPI using RPCBF generates the safest trajectories. The statistics of the safety performance of the controllers are shown in table I.
Figure 7: Trajectory Comparisons on AutoRally. SMPPI-RPCBF (in blue) leads to the tightest spread of states inside the track.
TABLE I: Collision & Crash Rate on AutoRally. MPPI causes the most collisions and crashes. SMPPI-HOCBF and SMPPI-PCBF improve safety, while SMPPI-RPCBF minimizes collisions & crashes.
@@
IV-D Comparison of Computation Times
Figure 8 shows the real-time feasibility of RPCBF-SF on the DI and F-16, highlighting how (component) computation times scale with increasing $T$ and $N$, evaluated on a laptop CPU.
Figure 8: Per-Timestep Computation Times for Double Integrator and F-16. Computation times for the total and individual components for varying $H$ ($N=64$) and NN ($H=50$), including mean and standard deviation across initial conditions and timesteps. The black box highlights the settings considered in the hardware experiments.
@@ V Hardware Experiments
We conduct hardware experiments on the Crazyflie platform to test the robustness of the proposed RPCBF-SF to real-world disturbances (see fig. 1).
Figure 9: Hardware Traces. RPCBF-SF maintains safety despite model errors, while PCBF-SF, assuming perfect dynamics, collides.
@@
We run the (R)PCBF-SFs with $\alpha=5$ for $6$ nominal trajectories; fig. 9 shows results for $3$ of them. The RPCBF-SF remains safe throughout, while the PCBF-SF always collides.
Figure 10: Safety for different $\boldsymbol{\alpha}$. Minimum distance to the obstacle over 66 random nominal trajectories. While PCBF is safe for small $\alpha$, it requires fine-tuning, whereas RPCBF is safe for all tested $\alpha$.
We next vary the choice of the class-$\kappa$ function $\alpha$ and plot the results in fig. 10. While the non-robust PCBF-SF does not collide with the obstacle when $\alpha$ is sufficiently small, this requires fine-tuning and is difficult to know beforehand. On the other hand, the RPCBF-SF is safe for all values of $\alpha$ we tested, allowing $\alpha$ to be used as a parameter that controls the behavior without also simultaneously affecting safety.
@@ VI Discussion and Conclusion
In this work, we proposed the Robust Policy Control Barrier Function (RPCBF), a method for constructing robust CBFs using the robust policy value function derived from rolling out a design policy. Subsequently, we introduced a real-time approximation that can be derived online, with conditions for its validity as a CBF. Simulation experiments demonstrate that a safety filter constructed using the RPCBF yields improved safety and more accurate estimation of the robust control-invariant set compared to existing methods. Hardware experiments on a quadcopter highlight the importance of accounting for model errors to ensure safety.
Future work will focus on analyzing the safety guarantees of the RPCBFs approximation, considering the finite-horizon, time-discretization, and sampling-based approach. Additionally, while the RPCBF acts as a CBF for any design policy if a long enough horizon is considered, conservativeness depends on the design policy. Deriving a policy to reduce conservativeness while maintaining infinite horizon guarantees is an important research direction. Moreover, extending our method to time-varying constraints and integrating real-time onboard perception are important directions for future work.